OpenStudy (anonymous):

The line including the two points (5, 3), (8, -2)

OpenStudy (triciaal):

First find the slope change in y/change in x use the values for one of the points and find the intercept b now write the eqn in the form y = mx + b

OpenStudy (anonymous):

5x -9x = -40?

OpenStudy (triciaal):

what is the change in y

OpenStudy (anonymous):

5x -9y = -40

OpenStudy (triciaal):

doing it now

OpenStudy (triciaal):

slope -5/3 -2 = -5/3 (8) + b b = 34/3 y = -5/3 x+ 34/3

OpenStudy (anonymous):

ok thank you

OpenStudy (triciaal):

do you know how to verify